OpenStudy (jacobhucakby1):

what is ratio of 12/2

OpenStudy (mathstudent55):

Treat 12/2 as a fraction and reduce it.

OpenStudy (mathstudent55):

What is the largest number that both 12 and 2 divisible by?

OpenStudy (will.h):

he's thinking give em a min

OpenStudy (johnny_richard2):

Divide each component by 2 to get 12/2 = 6 and 2/2 = 1 So the ratio 12:2 reduces to 6:1

OpenStudy (will.h):

there you go

OpenStudy (courtney233):

12 /2 dived by 2 = 6/1